# RedNote-Vibe: A Dataset for Capturing Temporal Dynamics of AI-Generated Text in Social Media

### v1

@@ 4 Experiments
Table 2: Comparison of PLAD with existing methods across detection tasks derived from RedNote-Vibe. All results are shown as percentages without the % symbol. Acc. denote accuracy. Precision, recall, and accuracy are computed as macro-averages.
@@
StyloAI 21.68 22.82 23.20 37.33 37.24 41.21 75.23 72.55 77.90
Ullah et al. 23.83 25.29 26.01 41.05 41.04 46.41 75.91 73.83 78.59
Binoculars 21.13 22.67 24.05 40.48 39.14 43.72 72.07 74.17 77.89
@@
BERT-base 32.97 32.71 33.76 44.38 42.41 50.47 85.91 88.29 88.24
ALBERT-base 32.26 31.74 31.65 41.98 40.61 44.34 83.59 79.63 84.58
RoBERTa-base 35.20 31.78 32.21 38.22 40.20 47.64 87.85 88.88 89.52
Sniffer 30.63 30.77 30.04 44.39 42.88 45.65 82.55 81.30 84.45
POGER 31.28 30.16 31.88 45.41 44.68 47.16 79.17 82.03 84.89
LLM-Idiosyncrasies 32.91 33.16 32.31 49.81 45.71 49.96 88.09 90.15 89.07
PLAD Framework with Different Classifiers (ours)
PLAD w/ GBC 30.61 31.27 31.79 49.41 48.00 52.83 86.16 85.32 87.63
PLAD w/ XGBoost 32.11 33.51 34.04 50.73 48.77 53.30 86.45 85.31 87.79
PLAD w/ CatBoost 35.87 36.45 36.94 50.06 47.34 51.89 88.70 87.28 89.62
4.1 AIGT Detection
@@
Experiment Setup. To evaluate the performance of our proposed PLAD, we compare it with two categories of approaches. (1) Classical statistics-based methods: StyloAI (Opara, 2024), Binoculars (Hans et al., 2024) and the method of Ullah et al. (2024), which serve as representative feature-driven baselines for AIGT detection. (2) Model-based methods: This category covers strong baselines, including (i) fine-tuning pre-trained text classification models such as BERT-base (Devlin et al., 2019), RoBERTa-base (Liu et al., 2019), and ALBERT-base (Lan et al., 2019), which represent state-of-the-art paradigms (Gritsai et al., 2024; Li et al., 2024), and (ii) established AIGT detection pipelines such as Sniffer (Li et al., 2023), POGER (Shi et al., 2024), and LLM-Idiosyncrasies (Sun et al., 2025). We follow the original training protocols of all methods and fine-tune them on our dataset. For comparison, we evaluate our proposed PLAD framework with different classifiers, including CatBoost (Prokhorenkova et al., 2018), XGBoost (Chen & Guestrin, 2016), and Gradient Boosting Classifier (Friedman, 2001).
Results. Table 2 presents the evaluation results across three detection tasks. It can be seen that classifiers with our PLAD framework outperform existing methods on most metrics. For the most challenging model identification task, PLAD with CatBoost achieves the best overall performance. This represents a notable improvement over existing approaches. The consistent performance across all metrics suggests that the psycholinguistic features effectively capture distinctive patterns among different LLMs. For the provider identification task, PLAD with XGBoost demonstrates the strongest performance. While LLM-Idiosyncrasies shows competitive precision, our approach maintains more balanced performance across all metrics. For the AIGT classification task, the results show a more competitive landscape. LLM-Idiosyncrasies achieves the highest recall, while our PLAD framework with CatBoost attains the best precision and accuracy. RoBERTa-base also demonstrates strong performance. The overall performance trend across tasks reflects the challenge of capturing subtle stylistic differences between similar LLMs, highlighting the value of our approach in providing interpretable insights.
4.2 Zero-shot Experiment
Table 3: Zero-shot performance comparison on the Provider Identification (6-way) task. Acc. denotes accuracy on the testing set of the seen model.
Method Unseen Acc. 0-shot Recall
PLAD GPT-o3 46.46 56.34
BERT-base GPT-o3 42.42 25.35
PLAD Gemini-2.5 43.72 58.46
BERT-base Gemini-2.5 45.23 52.31
@@ 4.2 Zero-shot Experiment
Given the observation that model providers often share similar training methodologies, data sources, and architectures across different versions. Therefore, a model family inherits a specific style imprint (Spiliopoulou et al., 2025). In this experiment, we evaluate the generalizability of detection methods on unseen AI models. We design a zero-shot experiment by excluding the latest GPT-o3 and Gemini-2.5 from the training data while retaining other models from their respective providers (OpenAI and Google). This setup mimics the scenario that detection systems encounter newly released models that are not available during training. We compare our PLAD framework and fine-tuning BERT-base on the reduced dataset, then evaluate their accuracy on seen models and recall on unseen target models. As results are shown in Table 3, our PLAD framework significantly outperforms BERT in identifying unseen models, demonstrating that the psycholinguistic features extracted by PLAD can capture more robust and generalizable traces of model families. In the context of rapid iteration of LLMs, this detection capability of new models makes our framework more practical than model-based methods.
@@ 4.3 Ablation and Feature Importance Study
To understand how PLAD achieves its detection performance, we conduct an analysis combining a dimension-level ablation study and evaluation of feature importance using CatBoost classifier. Figure 3 and Table 4 present the feature importance and ablation study result. It reveals that the largest performance drop occurred when the Lexical Identity and Stylistic Signature features are removed, leading to a substantial decrease across all three tasks. This finding identifies the two most important features, Prosodic Rhythm Consistency (17.6) and Type-Token Ratio (13.2), suggest that AI-generated text tends to exhibit smoother and more uniform rhythmic patterns, whereas human writing often contains irregularities caused by cognitive processing. The third most important feature, Imperfection (5.2), also from this dimension, which detects the absence of human-like hesitations, self-corrections, and other disfluencies that LLMs are optimized to avoid.
In addition, the analysis highlights the significance of Emotional and Social Grounding. Features from this category, such as Interactive and Dialogic Stance (4.8) and Emoji Density (3.0), rank highly in importance. This demonstrates PLAD’s ability to distinguish the subtle markers of human social awareness and interaction from the engagement patterns produced by AI. In summary, PLAD identifies AI content by recognizing its too perfect, too uniform, and lack of social characteristics. We also analyze the differences between human and AI text in Appendix.
Figure 4: Temporal evolution of AI content proportion in exploration set, where visualizations are smoothed using 14/30-day rolling averages.
@@ 5 Analysis
In this section, we use PLAD framework to analyze the exploration set in the RedNote Engagement Dataset, uncovering the impact of AI-generated content on user behavior and interaction trends in real-world data with temporal dynamics.
5.1 Temporal Dynamics of AI Content
Figure 4 presents the temporal track of AI-generated content proportion over the past 600 days. Despite considerable short-term noise, the overall linear trend ($slope=0.012$ per day, $0.355$ per month) clearly demonstrates a steady rise in the adoption of AI-authored posts.
In the initial phase, the proportion of AI content remained relatively modest and exhibited substantial volatility. However, from mid-2024 onward, the trend reveals a more persistent upward shift, indicating that AI tools are becoming a significant part of social media activity. It is also notable that the track does not follow a simple monotonic increase. Several plateaus and temporary declines can be observed (e.g., mid-2024 and mid-2025). These inflection points coincide with platform-level governance measures on AI-generated content, as recorded in the news during the same period. Overall, the findings indicate that while AI adoption in social media continues to grow steadily in the long term, its short-term dynamics remain subject to both platform governance and stochastic factors.
5.2 Engagement Metrics of AI vs. Human Posts
Figure 5: Engagement metrics comparison between Human-authored and AI-generated posts across 6 major topical domains. Human content achieves higher metrics than AI content in most domains.
To investigate engagement disparities, we compare metrics for human-authored and AI-generated posts across the 6 most data-rich domains. Given the heavy-tailed nature of interaction data, we apply a base-10 logarithmic transformation to normalize the distributions. Figure 5 presents these comparisons using boxplots, with blue indicating human content and orange indicating AI content.
@@ 5.1 Temporal Dynamics of AI Content
The analysis reveals that human-authored posts consistently outperform their AI counterparts, achieving higher median interaction counts and a greater propensity for high engagement (i.e., longer upper whiskers). This gap is particularly visible in domains requiring nuanced personal experience and emotional resonance, such as Travel, Career, and Relationships. This suggests that current AI content is significantly weaker than human content at establishing emotional resonance.
Furthermore, compared to human content, the upper whiskers of AI content are shorter. This indicates that the interaction for AI content is more concentrated with less variance. In other words, AI content tends to be homogeneous and is less likely to produce the exceptionally high-engagement posts. Conversely, human content’s longer upper whisker demonstrates its capacity to produce breakout posts that capture the collective imagination. This virality is often driven by novelty, raw emotional expression, or a unique personal story that breaks a predictable pattern.
Looking into the variations between interaction types, which reveals the cognitive investment from users. For instance, a like is a passive and lightweight social signal, whereas a comment requires active and conscious engagement. Human content shows a steep decline from likes to comments across most domains, while AI content maintains a relatively flatter curve. The Fashion domain presents an exception, showing comparable engagement patterns. These suggest that AI-generated content, while receiving lower absolute engagement, tends to provoke more discussion relatively.
@@ 5.3 Author-Level AI Usage and Engagement
In addition to content-level analysis, we conduct an author-centric study to understand how different AI usage behaviors correlate with engagement. We select authors with a minimum of four posts to ensure sufficient data for usage calculation, resulting in a total of 829 authors. Based on the proportion of AI-generated content they produced, we categorize these authors into three tiers, which consist of 68.7% Traditional Authors, who compose manually without AI; 27.7% AI-reliant Authors, who fully rely on AI for posting; and a small but distinct group of 3.5% of AI-augmented Authors, who combine human- and AI-generated content to create posts.
As the results illustrated in Figure 6, the Traditional Author and AI-reliant Author groups have similar trends to the analysis in the previous section. However, the AI-augmented Authors group consistently achieves higher engagement across all three metrics. This finding suggests that a balanced, strategic integration of AI tools outperforms approaches that rely exclusively on either human or AI-driven creation. The most successful authors are leveraging AI while applying human creativity to create more engaging and popular content.
6 Related Work
Social media datasets have been developed for AIGC detection, including TweepFake (Fagni et al., 2021) for early Twitter content, SAID (Cui et al., 2023) for modern LLM detection, and ElectionRumors2022 (Schafer et al., 2024) for election-related content analysis. Chinese social media datasets focus primarily on Sina Weibo for tasks like NER (Peng & Dredze, 2015) and fake news detection (Yang et al., 2021), but lack coverage of RedNote platform and temporal dynamics.
AIGC detection methods span three categories: watermarking techniques that embed imperceptible marks during generation (Kirchenbauer et al., 2023; Liu et al., 2024), classifier-based approaches using fine-tuned transformers like BERT and RoBERTa (Hu et al., 2023; Huang et al., 2024), and statistical methods that establish discrimination thresholds (Mitchell et al., 2023; Zhang et al., 2024). Multi-class detection frameworks like Sniffer (Li et al., 2023) and LLM-Idiosyncrasies (Sun et al., 2025) leverage LLM embeddings for classification. However, existing statistical methods struggle with real-world accuracy (Qiu et al., 2024; Sadasivan et al., 2023), highlighting the need for more robust and interpretable approaches.
7 Conclusion
In this work, we introduce the RedNote-Vibe, the first comprehensive social media AIGC detection dataset that captures temporal dynamics and engagement patterns. To address the limitations of existing AIGC detection methods in social media contexts, we propose the PsychoLinguistic AIGT Detection Framework (PLAD), which adopts psychological and linguistic theories to achieve both high accuracy and interpretability. Leveraging our dataset and detection method, our analysis reveals several important insights: (1) AI content adoption on social media platforms shows steady growth but is sensitive to platform governance policies; (2) Human-generated content consistently achieves higher engagement, particularly in domains requiring emotional resonance; (3) Authors who combine human creativity with AI assistance achieve the highest engagement levels, suggesting potential benefits of human-AI collaboration. The code and datasets will be publicly available, which provides a new resource for future social media research.
@@ 7 Conclusion
In the Appendix, we provide further analysis, including the differences between human-written text and AI-generated text from the perspective of the PLAD framework (Section A.1) and how these features affect the engagement metrics of posts (Section A.2). In addition, we analyze the correlation between features, reflecting the orthogonality of our framework (Section A.3).
Appendix A Further Analysis
A.1 Feature Statistics for Human and AI Post

### v2

@@ 4 Experiments
StyloAI (2024) 21.68 22.82 23.20 37.33 37.24 41.21 75.23 72.55 77.90
Ullah et al. (2024)) 23.83 25.29 26.01 41.05 41.04 46.41 75.91 73.83 78.59
Binoculars (2024) 21.13 22.67 24.05 40.48 39.14 43.72 72.07 74.17 77.89
@@
FT. BERT (2019) 32.97 32.71 33.76 44.38 42.41 50.47 85.91 88.29 88.24
FT. ALBERT (2019) 32.26 31.74 31.65 41.98 40.61 44.34 83.59 79.63 84.58
FT. RoBERTa (2019) 35.20 31.78 32.21 38.22 40.20 47.64 87.85 88.88 89.52
Sniffer (2023) 30.63 30.77 30.04 44.39 42.88 45.65 82.55 81.30 84.45
POGER (2024) 31.28 30.16 31.88 45.41 44.68 47.16 79.17 82.03 84.89
LLM-Idiosyncrasies (2025a) 32.91 33.16 32.31 49.81 45.71 49.96 88.09 90.15 89.07
PLAD Framework with Different Classifiers
PLAD-XGBoost 32.11 33.51 34.04 50.73 48.77 53.30 86.45 85.31 87.79
PLAD-CatBoost 35.87 36.45 36.94 50.06 47.34 51.89 88.70 87.28 89.62
Table 4: Comparison of PLAD with existing methods across detection tasks derived from RedNote-Vibe. FT. denote fine-tuning. Precision, recall, and accuracy are computed as macro-averages.
4.1 Benchmarking Detectors on RedNote-Vibe
@@
Experiment Setup. We compare AIGT detectors on our proposed tasks, with three categories of approaches. (1) Statistics-based methods: StyloAI (Opara, 2024), Binoculars (Hans et al., 2024) and the method of Ullah et al. (2024), which serve as representative feature-driven baselines for AIGT detection. (2) Model-based methods: This category covers strong baselines, including (i) fine-tuning pre-trained text classification models such as BERT-base (Devlin et al., 2019), RoBERTa-base (Liu et al., 2019), and ALBERT-base (Lan et al., 2019), which represent strong baselines (Gritsai et al., 2024; Li et al., 2024), and (ii) state-of-the-art AIGT detection pipelines such as Sniffer (Li et al., 2023), POGER (Shi et al., 2024), and LLM-Idiosyncrasies (Sun et al., 2025a). (3) PLAD: For comparison, we evaluate our proposed PLAD framework with different classifiers, i.e., CatBoost (Prokhorenkova et al., 2018) and XGBoost (Chen and Guestrin, 2016). We follow the original training protocols of all methods and implement them on our dataset. All methods share the same dataset and are split into training, validation, and testing sets in an 8:1:1 ratio.
Results. Table 4 summarizes the performance across three tasks. For AIGT classification, competitive results are observed, with LLM-Idiosyncrasies achieving the highest recall while PLAD-CatBoost attains the best precision and accuracy. In the more granular provider and model identification tasks, PLAD consistently outperforms baselines. These results suggest that our psycholinguistic features effectively capture the subtle stylistic signatures of different LLMs, offering both superior detection capability and interpretability.
4.2 Generalization to Unseen Models
Method Seen Acc. Unseen Recall
GPT-o3 PLAD 46.46 56.34
BERT-base 42.42 25.35
ALBERT-base 42.73 28.31
LLM-Idiosyncrasies 44.34 22.54
Gemini-2.5 PLAD 43.72 58.46
BERT-base 45.23 52.31
ALBERT-base 42.61 50.00
LLM-Idiosyncrasies 41.71 53.85
Table 5: Recall on unseen target models (GPT-o3 and Gemini-2.5) in Provider Identification (6-way) task.
@@ 4.2 Generalization to Unseen Models
In this experiment, we evaluate the generalizability of detection methods on unseen AI models. Given the observation that model providers often share similar training methodologies, data sources, and architectures across different versions. Therefore, a model family inherits a specific style imprint (Spiliopoulou et al., 2025). We design a zero-shot experiment by excluding the latest GPT-o3 and Gemini-2.5 from the training data while retaining other models from their respective providers (OpenAI and Google). This setup mimics the scenario that detection systems encounter newly released models that are not available during training. We set up experiments on the reduced dataset, then evaluate their accuracy on seen models and recall on unseen target models. As results are shown in Table 5, our PLAD framework significantly outperforms other approaches in identifying unseen models, demonstrating that the psycholinguistic features extracted by PLAD can capture more robust and generalizable traces of model families. In the context of rapid iteration of LLMs, this detection capability of new models makes our framework more practical than model-based methods.
@@ 4.3 Ablation and Feature Importance Study
Figure 3: Engagement metrics comparison between Human-authored and AI-generated posts across 6 topical domains.
To understand how PLAD achieves its detection performance, we conduct an analysis combining a dimension-level ablation study and evaluation of feature importance using CatBoost classifier. Table 6 present the ablation study result. It reveals that the largest performance drop occurred when the Lexical Identity and Stylistic Signature features are removed, leading to a substantial decrease across all three tasks. We further identify the two most important features, Prosodic Rhythm Consistency and Type-Token Ratio, suggest that AI-generated text tends to exhibit smoother and more uniform rhythmic patterns, whereas human writing often contains irregularities caused by cognitive processing. The third most important feature, Imperfection, also from this dimension, which detects the absence of human-like hesitations, self-corrections, and other disfluencies that LLMs are optimized to avoid. Detailed feature importance and the differences between human and AI text analysis are presented in Appendix B.
@@ 5 Analysis
In the previous experiments, we demonstrated that our proposed PLAD framework achieves state-of-the-art performance in detecting AIGT on the RedNote dataset, outperforming existing black-box baselines. We now apply PLAD to the real-world posts from the post-LLM era to investigate the temporal evolution of AIGT and its impact on social media engagement.
Considering that the exploration set (90,000 post-LLM posts) is inherently unlabeled, as only the authors themselves know whether AI was used, and human annotation is also infeasible. Therefore, we use pre-LLM posts to evaluate model reliability. We adopt binary AIGT classification and evaluate the model on pre-LLM validation set, which consists of posts written by humans and independent from the exploration set used in downstream analyses. To prioritize precision and minimize false positives, we set a conservative threshold of 0.9 based on validation performance, achieving a false positive rate (FPR) of 4.9%. Note that this threshold is determined solely from the labeled validation set before any analysis is conducted. After that, we apply the model with the pre-determined threshold to the naturally unlabeled exploration set for subsequent analyses. This process follows a validation-then-deployment pipeline and is consistent with previous AIGT analysis studies (Sun et al., 2025b; Liang et al., 2025).
Figure 4: Temporal evolution of AIGT percentage across different domains (2021-2025).
5.1 Temporal Evolution of AI Content
Figure 4 illustrates the monthly percentage of detected AI content across different domains. Since the emergence of LLMs, the proportion of AI-generated content has exhibited a steady growth across all domains. This trend indicates that AI tools are increasingly utilized for lifestyle sharing. This finding aligns with recent studies on other datasets Sun et al. (2025b); Liang et al. (2025), reinforcing the observation that the adoption of generative AI is a widespread phenomenon.
5.2 Metrics of AI vs. Human Posts
@@ 5.1 Temporal Evolution of AI Content
To investigate the impact of content source on interaction, we compare engagement metrics across six representative domains in Figure 3. It reveals that Human content exhibits superior emotional resonance. Human-authored posts consistently outperform AI-generated ones in median engagement. This gap is particularly visible in domains requiring personal experience and emotional resonance, such as Travel, Career, and Relationships.
Compared to human content, the interaction distribution for AI posts is more concentrated, with notably fewer high-engagement outliers. This indicates that while AI can generate grammatically correct text, it lacks the novelty and high-arousal qualities necessary to create breakout content. Looking into the variations between interaction types reveals the cognitive investment from users. For instance, a like is a passive and lightweight social signal, whereas a comment requires active and conscious engagement. Interestingly, the human-AI gap in comments is smaller than in likes. This suggests that while AI content receives lower overall engagement, the performance gap narrows for interactions requiring higher cognitive investment.
@@ 5.3 Author-Level AI Usage and Engagement
In addition to content-level analysis, we conduct an author-centric study to understand how different AI usage strategies correlate with user interaction behavior. We select authors with at least four posts in our dataset to estimate their AI usage, resulting in 829 authors. We compute each author’s AI ratio by their posts classified as AIGT by PLAD. Based on this, we categorize authors into three groups: (1) Traditional Authors (AI ratio = 0%), who compose entirely without AI assistance (570 authors, 68.7%); (2) AI-reliant Authors (AI ratio = 100%), who fully rely on AI for all their posts (230 authors, 27.7%); and (3) AI-augmented Authors (0% << AI ratio << 100%), who strategically combine human and AI-generated content (29 authors, 3.5%). For each author group, we present their engagement metrics in Figure 5.
The results of Traditional Author and AI-reliant Author groups are consistent with the findings in the previous section. However, the AI-augmented Authors group achieves relatively higher engagement. We hypothesize that these authors leverage AI for efficiency while applying human judgment to refine the content quality. This might suggest that a small group of users who strategically utilize AI tools can achieve higher engagement outcomes, outperforming both those who avoid AI entirely and those who rely on it exclusively.
6 Discussion and Conclusion
Rethinking AIGT Detection in the Wild. Existing detection methods often report high accuracy (e.g., $98\%$). However, we argue that this performance is inflated by the stylistic gap between naive LLM outputs and human writing. In this work, we optimized our generation pipeline to produce content more indistinguishable from human posts. It makes RedNote-Vibe more challenging, reflecting real-world adversarial conditions.
Longitudinal Interactions Between Human and AI Content. As AIGT becomes increasingly prevalent on social media, users are exposed to both human- and AI-generated content with growing frequency. LLMs are trained on human-produced corpora, while their outputs may subsequently influence human writing patterns. Our dataset provides a resource for investigating these temporal dynamics, which we leave for future work.
Conclusion. In this paper, we introduce a novel dataset that fills the gap in Chinese lifestyle social media resources. We also propose an AIGT detection method as the baseline for our dataset and tasks. Our exploratory analysis reveals some AIGT and user behavior patterns of social media. We emphasize that our core contribution lies in the dataset itself, we believe it provides a valuable data source for future AI/human behavior research.
Ethical Statement
All data utilized in this study originates from publicly accessible posts on the RedNote platform. We do not collect any private information. For the author-centric study, we hashed the usernames for grouping, and we only performed statistical analysis, not on individuals.
Regarding the AI-generated variants, they are produced using commercial LLM APIs with the sole objective of constructing a robust detection benchmark. We explicitly prohibit the misuse of this dataset for generating misleading content or spam.
Limitations
While our work provides comprehensive insights into AIGT on social media, we acknowledge several limitations that future work should address:
Language and Platform Specificity. Our study focuses exclusively on RedNote (Xiaohongshu), a platform with dominant of Chinese-language users. Consequently, the linguistic features and engagement patterns observed here may not directly generalize to other platforms. Furthermore, although our PLAD framework is designed based on general psycholinguistic principles and is theoretically language-agnostic, we have not empirically verified its efficacy across other languages in this study.
Limited Coverage of Generative Models. The AIGT variants in our dataset are generated using a subset of LLMs available up to mid-2025. Given the rapid iteration of model capabilities (e.g., the emergence of reasoning-heavy models), the generated vibe may continue to evolve.
Reliance on Proxy LLMs. Our feature extraction pipeline relies on a proxy LLM to quantify psychological dimensions. Biases present in the proxy LLM could propagate into the feature set. However, these biases are unlikely to significantly impact the analysis results. In the Appendix C.3, we provide a robustness evaluation, demonstrating that our method outperforms existing approaches in terms of correlation with human judgments.
Cost of Inference. Compared to lightweight end-to-end classifiers (e.g., RoBERTa-based detectors), PLAD involves a multi-step process of feature extraction via LLMs, which incurs higher computational and financial costs. This may limit its real-time application in high-throughput industrial scenarios without further optimization.
@@ 6 Discussion and Conclusion
In the Appendix, we provide related work (Section A) and further analysis, including the differences between human-written text and AI-generated text from the perspective of the PLAD framework (Section B.1) and how these features affect the engagement metrics of posts (Section B.2). In addition, we analyze the correlation between features, reflecting the orthogonality of our framework (Section B.3). Finally, we introduce details of our AI samples generation (Section C.1), feature extraction method (Section C.2) and the validation of our feature quality (Section C.3).
Appendix A Related Work
Social media datasets have been developed for AIGC detection, including TweepFake (Fagni et al., 2021) for early Twitter content, SAID (Cui et al., 2023) for modern LLM detection, and ElectionRumors2022 (Schafer et al., 2024) for election-related content analysis. Chinese social media datasets focus primarily on Sina Weibo for tasks like NER (Peng and Dredze, 2015) and fake news detection (Yang et al., 2021), but lack coverage of RedNote platform and temporal dynamics.
AIGC detection methods span three categories: watermarking techniques that embed imperceptible marks during generation (Kirchenbauer et al., 2023; Liu et al., 2024), classifier-based approaches using fine-tuned transformers like BERT and RoBERTa (Hu et al., 2023; Huang et al., 2024), and statistical methods that establish discrimination thresholds (Mitchell et al., 2023; Zhang et al., 2024). Multi-class detection frameworks like Sniffer (Li et al., 2023) and LLM-Idiosyncrasies (Sun et al., 2025a) leverage LLM embeddings for classification. However, existing statistical methods struggle with real-world accuracy (Qiu et al., 2024; Sadasivan et al., 2023), highlighting the need for more robust and interpretable approaches.
Appendix B Further Analysis
B.1 Feature Statistics for Human and AI Post
Figure 6: The top-15 important feature scores.